OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=-2x+7 y=1/3x

OpenStudy (anonymous):

so what are we finding out?

OpenStudy (anonymous):

The solution to the system so an ordered pair

OpenStudy (anonymous):

@YoungStudier

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Lexi_Loves

OpenStudy (lexi_loves):

ohh ok so do you want to graph it?

OpenStudy (anonymous):

Ya I need to find the ordered pairs

OpenStudy (lexi_loves):

Are you having trouble working the problem out?

OpenStudy (anonymous):

Ya I don't really understand it

OpenStudy (lexi_loves):

ok.

OpenStudy (anonymous):

So can you help me?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (lexi_loves):

I dont want to give you he wrong answer!

OpenStudy (anonymous):

I know

OpenStudy (lexi_loves):

(1,7) idk..

OpenStudy (anonymous):

ok I will try it

OpenStudy (lexi_loves):

????

OpenStudy (anonymous):

It was (3,1) @Lexi_Loves

OpenStudy (lexi_loves):

really how? im sorry

OpenStudy (luigi0210):

You could of found the order pair by sub as well: \(\Large y=\frac{1}{3} x \) \(\Large y=-2x+7 \) ---------------------------- \(\Large \frac{1}{3} x= -2x+7 \) ---------------------------- \(\Large x=-6x+21 \) ---------------------------- \(\Large 7x=21 \) ---------------------------- \(\Large x=3 \) --------------------------- \(\Large y=\frac{1}{3} (\color{red}{3}) \) \(\Large y=\color{red}{1} \) ----------------------------------- (x, y) (3,1) Sorry I was late

OpenStudy (anonymous):

It's ok @Lexi_Loves I still passed

OpenStudy (anonymous):

And thank you @Luigi0210 you just taught me how to do that